Direction: Solve the problem below and construct probability distribution and make a histogram.




Suppose three cellphones are tested at random. We want to find out the number of defective cellphones that occur. Thus, to each outcome in the sample space we shall assign a value.

1
Expert's answer
2022-03-04T12:21:30-0500

Let "D" represent the defective and "N" represent non-defective cellphone. Let "X" be the random variable representing the number of defective cell phones.

Sample space:


"S=\{NNN,NND, NDN,DNN,"


"NDD, DND, DDN,DDD\}"

The possible values of the random variable "X" are "0, 1, 2, 3."

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Possible \ Outcomes & X \\ \hline
 NNN & 0\\
 \hdashline
 NND & 1 \\
 \hdashline
 NDN & 1 \\
 \hdashline
 DNN & 1 \\
 \hdashline
 NDD & 2 \\
 \hdashline
 DND & 2 \\
 \hdashline
 DDN & 2 \\
 \hdashline
 DDD & 3 \\
 \hdashline
 
\end{array}"


We will assume that the probability of getting heads and tails is the same:


"p = q =1\/2"



Construct the probability distribution of the random variable


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 0 & 1 & 2 & 3 \\ \hline
 p(x) & 1/8 & 3/8 & 3/8 & 1/8 
\end{array}"

or


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 0 & 1 & 2 & 3 \\ \hline
 p(x) & 0.125 & 0.375 & 0.375 & 0.125 
\end{array}"